OpenStudy (anonymous):

A 1.50 x 10^3 kg car leaves a parking lot. Thirty seconds later it is moving along a highway at 72km/hr. a. What is the car's change in momentum b. What average force does the motor produce to bring about this change in momentum.

OpenStudy (anonymous):

whats p? and what problem?

OpenStudy (anonymous):

so isn't it the change in momentum we are looking for?

OpenStudy (theeric):

Isn't it \(p=mv\) ?

OpenStudy (anonymous):

isn't it change in p = mvf-mvi?

OpenStudy (loser66):

oh yea, let theEric help us. I am not good at Physics

OpenStudy (theeric):

@jacobsmith222 Yeah, the change would be like \(p_\text{72km/h}-p_\text{parked}=\Delta p\)

OpenStudy (anonymous):

then what is p=mv?

OpenStudy (theeric):

That is the definition of momentum. Thanks @Loser66! You got kinetic energy there, hehe... :)

OpenStudy (theeric):

So \(p\) is momentum. And \(m\) is the mass. And \(v\) is the velocity.

OpenStudy (anonymous):

okay that makes sense. Would it be initial or final velocity for v in the definition

OpenStudy (theeric):

So you see that it's a simple equation! And not hard to remember! If you ever got hit by something, you want it to have less momentum. Lots of momentum: football player, 200 pounds, running full speed. Not much momentum: a fly buzzing around, lazily. \(p=mv\)

OpenStudy (anonymous):

okay so how would we solve this problem?

OpenStudy (theeric):

When you look at the change, it's always \(\text{final - initial}\). And here's a trick for substituting. When you look at \(p_\text{parked}\) for example, you replaced it with \(m_\text{parked}v_\text{parked}\). Of course, \(m\) is always the same, so you can just say \(mv_\text{parked}\).

OpenStudy (anonymous):

Okay so is it going to be 1.5x10^3(72)-1.5x10^3(0)

OpenStudy (anonymous):

also for the problem yesterday about finding the magnitude of the impulse, is it NS and change of p=FxT

OpenStudy (theeric):

Yeah, that's the calculation for part (a)!

OpenStudy (anonymous):

and final units is kg/km/hr

OpenStudy (theeric):

actually, kg * km / hr since its mass * velocity. :)

OpenStudy (anonymous):

okay and also for the problem yesterday about finding the magnitude of the impulse, is it NS and change of p=FxT

OpenStudy (theeric):

Newton seconds for impulse? I think so!

OpenStudy (anonymous):

okay and how would we achive b?

OpenStudy (theeric):

Was I in a problem with impulse yesterday? Hehe... I don't remember...

OpenStudy (anonymous):

It was A force of 30 N is applied to a hockey puck for .8s. Find the magnitude of the impulse

OpenStudy (theeric):

So, impulse if often \(J\) - is that what you're using?

OpenStudy (anonymous):

p

OpenStudy (theeric):

Ah, see, I use \(p\) for momentum and \(J\) for impulse. If you use \(p\) for impulse, what do you use for momentum? I would do \(J=\bar F \times t\).

OpenStudy (theeric):

But I think there is a more helpful equation.

OpenStudy (anonymous):

p=fxt

OpenStudy (theeric):

\(J=\dfrac{\Delta p}{\Delta t}\) by definition, right?

OpenStudy (theeric):

Scratch that - I want to use your variables.

OpenStudy (anonymous):

wait before we do this. How do I find b?

OpenStudy (theeric):

Haha, we can combine our two impulse equations... We can find impulse with my last equation, and we can solve for \(\bar F\) in your first equation. The bar just means "average." What variable do you use for momentum?

OpenStudy (anonymous):

p

OpenStudy (theeric):

And what do you use for impulse?

OpenStudy (anonymous):

idk. All we have is change in p =fxt

OpenStudy (theeric):

I made a mistake before, \(\bar F=\dfrac{\Delta p}{\Delta t}\), sorry! But that's how you would do part (b)... THAT's the definition that will help us!

OpenStudy (theeric):

And that, really, is the same as your equation. No impulse involved there! :)

OpenStudy (anonymous):

so what exactly would i plug in

OpenStudy (theeric):

\(\bar F=\dfrac{\Delta p}{\Delta t}\Longleftrightarrow \Delta p=\bar F\ \Delta t\)

OpenStudy (anonymous):

could i just do p=ma and find a and plug it in?

OpenStudy (theeric):

\(F=ma\), not \(p=ma\)! And it will take some work to find \(a\), and you don't have to! But that is one way. But give my equation a shot. Force is the change in momentum over a change in time. You know both the change in momentum (you calculated it) and the change in time (it's given in the problem).

OpenStudy (anonymous):

108000/30

OpenStudy (anonymous):

3600

OpenStudy (theeric):

Mind the units! :) Sorry for my delay...

OpenStudy (theeric):

Notice the momentum is in kg km/hr and your time, 30, is in seconds. So your 3600 is 3600 kg*km/hr/s, which isn't quite Newtons! It actually isn't a unit of force, without the time unit being the same, I think.

OpenStudy (anonymous):

I got 3600kgm/s^2

OpenStudy (theeric):

Oh, haha, sorry! Let me check. Those would be the right units, sorry! :)

OpenStudy (anonymous):

also cna you help me out with 1. A force of 8.0 N acts on a 2.0 kg mass for 5.0s a. what is the change in momentum of the mass b. what is the change in the velocity of the mass 2. A force of 6.00 N acts on a 3.00 kg object for 10.0 s a. What is the objects change in momentum b. What is its change in velocity

OpenStudy (theeric):

I suggest you ask that in another post! That way others can see and help too. I'll look at it in between my distractions!

OpenStudy (anonymous):

thanks!!

OpenStudy (anonymous):

Will they be different since its the objects change in momentum and velocity and not the change in velocity and momentum of the mass?

OpenStudy (theeric):

Change in momentum is 108,000kg*km/hr. Change in time is 30s. So average force is \(\dfrac{108,000\text{ [kg*km/hr]}}{30\text{ [s]}}=3,600\text{ kg*km/hr/s}\) And since... \(\mathrm{ 1km=1000m}\) \(\mathrm{1hr=3600s}\) we see by substitution that... \(\mathrm{3,600 kg*km/hr/s=3,600kg*1,000m/3600s/s=1,000 kg*m/s^2=1,000N}\) I think those are the right units.

OpenStudy (anonymous):

I apologize but where did 108000 come from

OpenStudy (theeric):

That is the change in momentum! You actually calculated it above! (see the attachment) \(\mathrm{1.50\times10^3kg\quad\times\quad72km/hr\quad-\quad1.50\times10^3kg\quad\times\quad0\\=108,000\ kg*km/hr}\)

OpenStudy (anonymous):

oh sorry I meant for my new problem

OpenStudy (theeric):

Post a new one! There will be much less clutter that way..

OpenStudy (anonymous):

2. A force of 6.00 N acts on a 3.00 kg object for 10.0 s a. What is the objects change in momentum b. What is its change in velocity